What is the oxidation number of chromium in Fe2(Cr207)3? Multiple Choice Ο Ο Ο Ο
Answer #1

What is oxidation no. of fea (02) ? 1 h fea (CRO) Fe3+ + cadre CY, 072- 22 - 14 =- 2 2n = 14-2=12 a = +6 onidation state of C